UT 09-Jan-2020

Night Center 00:26

Plan

-2nd half Eng/D-time
  • AGW 1/3 Transform Data collection ~ 1hr LUCI-Bino
    • Either confirmatory or new set depending on analysis from D. Miller
    • New transforms ready to be implemented. Stone E avail midnight-3am; Stone I 3am to end of night.
    • Barry's instructions for implementing a transform:
ssh telescope@obs1
ssh oac@oac
cd /lbt/oacontrol/current/etc/
file is oacontrol.conf organized by AGW#.

See the related instructions in BGTechProcedures:
 Issue the AGw command to implement the new parameters in the UMAC: rdwrconfig -u [#] NOTE: You must stop and restart the appropriate GCS subsystem for the changes to take effect (although maybe not these particular changes?). 

edit file in the AGW4 section,
rdwrconfig -u4
then stop and restart GCSR


PCSINSTRUMENT
ssh -X tcs@tcs1 (vis obs1 as telescope)
/home/telescope/Configuration/PCS/PCSInstrument.conf
  • D-targets
    • Conrad+LPL
      • Echeclus (1-hr open-shutter, sets to 30-deg elev at 2:30am)
        • Echeclus, then telluric. Solar Analog is end-of-night target.
      • C/2019A (above 30-deg between 10-12UT)
    • O. Kuhn QSO - LUCI (zJ and HK G200 1" slit) and MODS (dual grating 1" slit)
      • slit width may be modified according to seeing - LUCI to 0.75" and MODS to 0.8" or 1.2"
      • thin cirrus acceptable, seeing up to 1-1.5" acceptable
      • Q0352+123 - 1hour - zJ and HK - (visible until 2am)
      • Q0730+257 - 1.3 hour - zJ and HK - (visible until 6am)
      • Q1237+134 - ~1 hr - zJ and HK - (visible from 3am)
        • if end of night and there is time, then possibly switch to MODS to complete Q1237+134
    • B. Rothberg - Triple AGN - LUCI HK G200 1" slit
      • thin cirrus acceptable, seeing must be 1" or better on guider/wfs (goal is to separate close-together nuclei)
      • file locations:
        • RM507-5/4: /home/observer/share/Service_Observing/Eng_Jan_2020/Triple_AGNs_Rothberg_LD_2019B_007
        • RM507-1-3: /home/observer/Service_Observing/Eng_Jan_2020/Triple_AGNs_Rothberg_LD_2019B_007
        • OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/Triple_AGNs_Rothberg_LD_2019B_007
      • each requires 2 acquisitions to get all 3 nuclei
      • SDSS J080535.36+524553.2 - hits 30 deg at 6am - 1.5 hours
      • SDSS J083803.66+505509.5 - hits 30 deg elevation at 7am (visible through twilight) - 1.5 hours
    • Backup/Filler
      • ICRF3 - any conditions (~20min per)
        • Scripts are located in:
          • /home/observer/share/Service_Observing/Eng_Jan_2020/Backup/ICRF3_Rothberg_LD_2019B_006 - Rm507-5/4
          • /home/observer/Service_Observing/Eng_Jan_2020/backup/ICRF3_Rothberg_LD_2019B_006 - Rm507 1-3
          • /home/lbto/Service_Observing/Eng_Jan_2020/backup/ICRF3_Rothberg_LD_2019B_006 - OBS-N
        • 0426+273
        • 0614+570
        • 0854-108
        • 0509+293
        • 0509+406
        • 0513+276
        • 0535+001
        • 0600-000
        • 0559+422
        • 0600+063
        • 0637-128
        • 0641-135
        • 0702+612
        • 0804-055
        • 0845-068
      • Local Mergers - any conditions
        • /home/observer/share/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/Local_Mergers - RM507-5/4
        • /home/observer/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/Local_Mergers - RM507-1-3
        • /home/lbto/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/Local_Mergers - OBS-N
          • NGC 3656- 30min Bino, 1hr mono
          • NGC 2655- 30min Bino, 1hr mono (best between 2-5am due to telluric delta airmass)
          • Arp 156 -30min Bino, 1hr mono (best between 2-5am due to telluric delta airmass)
          • NGC 3597 (low elevation - never goes above airmass 1.7 - good if running out of targets) - 30min Bino, 1hr mono

Summary

We started with the transform verification, collecting data for ~25 stars in StoneE. The seeing was not so good to start with but then stabilized around 1.1-1.2" with a few excursions still, so we just took the full set. At the very end, last 2-3 stars, the seeing blew up.

By that time, it was almost 10 UT, the beginning of the 2-hr window for A/2019 C1 using bino LUCI G200 HK 1.5" slit: (LD-2019B-013, but note that solar analog has propid LD-2019B-012 as the script was generated from another LPL submission). We tried it - seeing was ~2-3" and we lost the guide star on SX for a while when it became swamped by high background, probably from a reflection of moonlight from something near the AGW1/L1 - unsure data will be useful, but managed to get the target+telluric and the solar analog at the end of the night, all at HK. At zJ we did not get any data.

After the non-sidereal observation, observed a z~2 quasar using bino LUCI G200 zJ and HK 1" slit (zJ and HK serially). (LD-2019B-011).

Hopefully the transform data confirm that the lfbg and rfbg transforms that Doug computed from the previous night's data are good. LUCIs worked well... there is some hint of a fuzzier spectrum on L2, but that really has to be looked into more.

During the transform data collection, on several occasions, the guide star on the right side was seen to oscillate about the hotspot, and the guider image appeared to be taken out of sync, i.e. while the probe was moving. Steve noted that this probably has to do with taking transform data; we did not see this behavior during the science observations that followed. We noted this also on the previous night.

The RTD makes looking at LUCI data in real-time challenging: it displays the last image and then quickly snaps back to the penultimate one. Then it changes the color/stretch. As a result, after each image is displayed, one has to click multiple times to get the last image back and adjust stretch. This is worse for the transform data collected through the N30 Field Stop since the contrast over the region displayed is very high. Then, finally, when one clicks on the image from one LUCI, the crosshairs and panning on the Aladin display for the image from the other LUCI changes. This behavior is on both 32 and 64-bit machines.

Log

06:54: ssh'd as oac@oac, cp oacontrol.conf oacontrol.bak.20200109

07:26 handover. Edited oacontrol.conf file to replace transforms (plus additional related parameters) with those which Doug supplied today (attached). Ran rdwrconfig -u1 and rdwrconfig -u3 and Steve restarted GCSL and GCSR.

08:00 Send on-axis preset to Stone_E GS=1. We have clouds and GS=1 is too faint. Try GS=3 instead.

08:03 On GCSR keep getting a timeout on WFS, but then it corrects self. Collimating - seeing FWHM ~ 1.0-1.2".
  • Steve resetting WFS controller to see if it helps with this issue.
  • Putting in N30 FS.
  • Steve homed probes.
  • Test exposures of 2.5 sec: L1 0005 and L2 0005 and at 3-sec about 10k counts. With clouds. I'll go for 4-sec to start with. But now the seeing is 2". We are pointing into the wind.
  • Giving it a try with list: Stone_Em135_0edit2.list which uses star 3.
    • WFS still not arriving within timeout. Had to control-C transform collect on the right hand side.

08:23 Right side still not WFSing. Left side did not find GS 70.

08:26 Steve is checking pointing. Stone E is 36 deg from the moon and we may be bothered by moonlit clouds. Stone I is further from the moon, but seeing may just deteriorate, so trying to stick with it. Seeing is 1.6" on average.

08:35 Giving it another try: Stone_Em135_0edit2.list.
  • Will start with L1 9 and L2 6
  • Note -- LUCI2 image quality looks a bit off but not consistently so.

L1 L2 GS comments
09 06 on-axis not so good seeing
10 07 background  
11 08 70 seeing ~1.2"  
12 09 9  
13 10 65  
14 11 21 just prior, collim reported L/R FWHM 1.47/1.2", AM=1.272
15 12 27 unsure whether GS27 was correctly found on both sides
16 13 85  
17 14 61 collimated just before, WFS FWHM 0.9"
18 15 45 45 is a faint GS
19 16 30  
20 17 36 collim just before, seeing ~1.1-1.2"
21 18 82  
22 19 80  
23 20 53 collim just before, seeing about same as before, problem with right side... not stable
transform_collect,exp=4,start=13 on right side, worked ok now but star was initially at upper left of guide window
24 21 4 gradient from probe
25 22 48  
25 23 84 what? took image 25 two times but different stars? that is what it says. On right, GS 84 not found - oscillated.
resending on right side after Steve restarted GCSR: transform_collect,exp=4,start=16
26 24 43  
27 25 20 gradient from probe
28 26 64 right side oscillating again... Steve restarting GCSR
29 27 89  
30 28 96  
31 29 118 on right side, seeing blew up and guide star hardly visible... out of sync so left was ok
32 xx 108 on right side, did not find guide star (09:31 UT) seeing blew up. Restarted with "2e" list which has 77 removed.
09:31 UT: transform_collect, exp=4.0, start=23 on right side
33 xx 77 left side: star out of range, skip it and start at star 92 (transform_collect,exp=4,start=25)
34 xx 92 star is OK - this is last star
started left side on new list, Stone_Em135_0edit4.list, while right side finishes
35 xx on-axis  
36 xx background  
37 xx 52  
38 xx 28  
39 xx 76  

On right side, gave up with original list and tried first "2e" and then just "2ee" to give it at least a completion. I think that by entering a new list and start=1, it went ahead and took the background images as if starting.

x 30 on-axis  
x 31 background  
x 32 82 new star added to "2ee"

Lists are in: ~/data/20200109/ and logs are in subdirectories called "lfbg" and "rfbg". Hope these data are helpful! Seeing did cooperate for the most part, except for the very beginning and the very end.

09:48 Done with transform data collection. Chiller high temp alarm had sounded and Steve waking Pat so they can have a look.

09:55 Going to a pointing star near A/2019A C1 (LPL) It is just about to rise above 30 deg limit. If this doesn't work, then quasars.
  • Seeing is about 2" here...
  • Trying - Edited guide star to move probe shadow further from source, just in case rates or offsets would cause vignetting.
  • No GS found on left side, but on right it is there. We just had corrected pointing - odd. Steve is spiraling on left side, but we see nothing!
  • Checking/correcting pointing again. Pointing was good.
  • Resent preset - star is found now on both sides. Seeing is about 2.5-3"
  • now (10:30) down to 1.6".
  • acquisition:
    • L1: 41,42,43(slit) and L2: 34,35,36(slit)
    • Object ID's at 42: (1016.27,577.12) and 35: (1036.31, 564.54)
    • Positioned on slit at L1 (1019.06, 580.44) and L2 (1025.62, 563.69)
    • 10:36: Sending offsets and taking confirmatory thru-slit images L1 44 and L2 37.

10:42 Think object is there in slit and starting the science. NB: Confirmed object is on the slit later by difference images L1 44-41 and L2 37-34. The RTD's quirks make looking at data in real-time pretty hard!!
  • We lost the guide star on the left side right away - it is buried in a high background. Moonlight is impacting left side.
  • Offset failed on left side. Managed to get right side pair + one exposure: L2 38,39 and 40 before offset failed.
  • Turned off left side "mirror", "instrument" and "readout" via observer execution panel, but again failed.
  • Offsetting by hand now.
    • Right side offset 0,0 will be L2 41 and right side offset 0,-10 will be L2 42.
11:00 Was going just to give up on object (seeing was poor, left side offset failed, and we had not requested the 25-deg elevation override which would have helped), but then the guide star emerged from the background on the left side.
  • Star is back on left side. Continuing both with offset 0,0. Will be L1 48 and L2 43.

11:15:57 A gust of wind must have knocked the telescope off-source. The preset was spontaneously canceled with the message: telescope off source for too-long. Came as a surprise since the guide/WFSing seemed good just before that. Apparently the wind was too strong. This ends the collection of data on the target.

11:16 Telluric for A/2019 C1 (LPL)
  • acquisition:
    • L1 54 star at (1019.06, 580.44) and L2 49 star at (1094.01, 652.15)
    • L1 55 slit (positioned star at 1034.53, 580.23)
    • L2 50 slit (positioned star at 1001.69, 563.94)
    • Sending offsets and taking confirmatory images: L1 56 and L2 51. These look good.
  • Taking series of spectra on telluric (AM=1.878). First, I mistakenly ran zJ script, but stopped after 1 exposure.
  • HK spectra on telluric are: L1 58-61 and L2 53-56.

11:42 Slew to Q1237+134 (BLRBH_Kuhn)
  • L1 acq images: 62,63 and 64 centered at 1058 on the slit
  • L2 acq images: 57,58 and 59 centered at 1033 on the slit
  • Confirmatory images: L1 65 and L2 60. Nudged to right by 0.1" for both sides and retook: L1 66, L2 61. Looks OK, hard to tell if there was a signficant change.
  • zJ start L1 67-70 and L2 62-65. Through-slit image does not look good on right hand side, although field image looks good. Perhaps object not quite as well centered on L2 as on L1.
  • More uncorrected flexure on L2. Also I have had to adjust FM4 position on L2 at the start of both 1/2 nights.

12:26: start HK on Q1237+134
  • L1 71-80 and L2 66-75

12:50: Slew to telluric for Q1237+134 (=HIP64248)
  • Guide star not found on left side.
  • acq images are: L1 81,82 and L2 76,77
  • zJ telluric L1 84-87 and L2 79-82
  • HK telluric L1 88-91 and L2 83-86

13:12: Solar Analog at HK
  • L1 92,93 and L2 87, 88
  • Confirmatory images (tried to put star around same Y as for comet & telluric). OK
  • HK spectra: L1 95-98 and L2 90-93
  • 13:30 repeating as it is the last target of the night and we can still guide, though well into twilight: L1 99-102 and L2 94-97. But on ~100 and 95 "spot determination failed" (i.e. it cannot collimate). Clouds are getting thicker... Tried to let this set of 4 finish, but clouds had other ideas... Guide star disappeared and last offset gave an error.

13:37 Closing up.

Cals for end of night:
  • 1.5" slit, G200 and HK (1.93 mic)
    • L1: 102-117 and L2: 097-112
  • 1" slit: G200 and HK (1.93 mic)
    • L1: 118-133 and L2: 113-128
  • 1" slit: G200 and zJ (1.17 mic)
    • L1: 134- and L2: 129-142

Now while telescope is at horizon for fills, taking some "darks" (blind filters in and 1" slit in FPU) for transform & for arcs:
  • 3x 4-sec darks (normal) for transform: L1: 148-150 and L2: 143-145
  • 3x20-sec darks for HK arcs: L1: 151-153 and L2: 146-148
  • 3x 60-sec for zJ arcs: L1: 154-156 and L2: 149-151
  • 3x 5-sec also for zJ arcs: L1 157-159 and L2: 152-154

-- BarryRothberg - 08 Jan 2020
Topic attachments
I Attachment Action Size Date Who Comment
lfbg.confconf lfbg.conf manage 630 bytes 09 Jan 2020 - 06:57 OlgaKuhn LFBG (AGW1) transform based on last night's data
rfbg.confconf rfbg.conf manage 630 bytes 09 Jan 2020 - 06:57 OlgaKuhn RFBG (AGw3) transform determined from last night's data
Topic revision: r9 - 09 Jan 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback